# Simultaneous extension of the polynomial Central Sets Theorem and the polynomial Milliken-Taylor theorem

### v1

@@ 2   Our proofs
Let $\alpha:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ be two functions such that $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$. For $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let $T_{G}=\{\sum^{m}_{i=1}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t):n\in\mathbb{N},\text{ each }G_{i}\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N}),G\subsetneq G_{1}\subsetneq G_{2}\cdots\subsetneq G_{n},\text{ and for each }i\in\{1,2,\ldots n\},f_{i}\in G_{i}\}$.
Now for $\alpha:\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ with $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$; define
@@
Then $Q(\alpha,H)\neq\emptyset$ and is a subgroup of $(\beta\mathbb{N},+)$. Furthermore,
$K(Q(\alpha,H))\subseteq(\mathcal{J},+).$
We show, by downward induction on $l\in\{1,2,\ldots,m\}$ that for each $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$,
@@
For $l=m$, we have $T_{G}\in p$ so that,
@@
$A=\left\{\sum_{i\in\beta_{l}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):\right.\\ \left.f_{i}\in G_{i},G\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l}<\beta_{l+1}<\cdots<\beta_{m}\right\}.$
We claim that,
$T_{G}\subseteq\{x\in\mathbb{N}:\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}\}:(x,x_{l+1},x_{l+2},\ldots,x_{m})\in A\}\in\bigotimes_{i=l+1}^{m}p_{i}.$
To this end let $b\in T_{G}$. Pick $G_{l}\in\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})$ with $G\subseteq G_{l}$ such that $b=\sum_{i\in\beta_{l}}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)$. Let $F=\cup_{i\in\beta_{l}}G_{i}$. Then
$\left\{\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+2}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):\right.\\ \left.f_{i}\in G_{i},F\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l+1}<\beta_{l+2}<\cdots<\beta_{m}\right\}\subseteq-b+A.$
So, $\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}:(b,x_{l+1},x_{l+2},\ldots,x_{m})\in A\}\in\bigotimes_{i=l+1}^{m}p_{i}$ as required.
Now we will prove Theorem 1.24, but before that let us recall one fact about ultrafilters from [13]. We will apply this result. For any discrete semigroup $(S,\cdot)$, and $p\in E\in\left(\beta S,\cdot\right)$, if $A\in p=p\cdot p$, then from [13, Lemma 4.14], the set $A^{\star}=\{x\in A:-x+A\in p\}\in p.$
@@
  2. ($2$)
    If $n\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\ldots\subsetneq G_{n}\subsetneq K$, then for each $j\in\{1,2,\ldots,m\}$, and $\beta_{1}<\beta_{2}<\dots<\beta_{j}\leq\{n\}$,
    $\sum_{i\in\beta_{j}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\in$
    $D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}$
@@
Then $M_{j}\subseteq D_{j}^{\star}$.
@@
$E=\bigcap^{m}_{j=1}\left\{D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}\right)\right)^{\star}:\right.\\ \left.\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T,\text{ and for each }i\in\{1,\dots,n\},f_{i}\in G_{i}\right\}$
$G=\bigcap_{\underset{j=1,\dots,m}{x\in M;}}-x+D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}\in P$
@@
The following theorem shows that the combinatorial consequence of our Corollary 1.25 is equivalent to a version of the polynomial extension of [13, Theorem 17.31]. The proof uses simple color induction arguments.
@@
    Let $r,m\in\mathbb{N}$, and let $[\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$. Then there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
      $\left\{\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right\}\subset C_{j}$
@@
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have for every $f_{i}\in G_{i},$ and $P\in T,$
      $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\in D_{j}$

### v3

@@ 2  Our Proofs
Let $\alpha:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ be two functions such that $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$. For $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let
$T_{G}(\alpha,H)=\left\{\sum_{i\in\beta}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right):n\in\mathbb{N},\,\beta\subseteq[1,n],G\subsetneq G_{1}\subsetneq G_{2}\cdots\subsetneq G_{n},\text{ and for each }i\in\{1,2,\ldots n\},f_{i}\in G_{i}\right\}.$
When $\alpha,H$ are fixed we write it $T_{G}$ instead of $T_{G}(\alpha,H).$
Now for $\alpha:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ with $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$; define
@@
It is easy to verify that $Q(\alpha,H)\neq\emptyset$ and is a subgroup of $(\beta\mathbb{N},+)$. Furthermore, $K(Q(\alpha,H))\subseteq(\mathcal{J},+).$ We show, by downward induction on $l\in\{1,2,\ldots,m\}$ that for each $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$,
@@
Let $p\in K\left(Q\left(\alpha,H\right)\right).$ For $l=m$, we have $T_{G}\in p$ so that,
@@
$A=\left\{\left(\sum_{i\in\beta_{l}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t))\right):\right.\\ \left.f_{i}\in G_{i},G\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l}<\beta_{l+1}<\cdots<\beta_{m}\right\}.$
We claim that
$T_{G}\subseteq\left\{x\in\mathbb{N}:\left\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}:(x,x_{l+1},x_{l+2},\ldots,x_{m})\in A\right\}\in\bigotimes_{i=l+1}^{m}p\right\}.$
To this end let $b\in T_{G}$. Pick $G_{l}\in\mathbb{P}_{f}(^{\mathbb{N}}{\mathbb{N}})$ with $G\subseteq G_{l}$ such that $b=\sum_{i\in\beta_{l}}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)$. Let $F=\bigcup_{i\in\beta_{l}}G_{i}$. Then
$\left\{\left(b,\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+2}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t))\right):\right.\\ \left.f_{i}\in G_{i},F\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l+1}<\beta_{l+2}<\cdots<\beta_{m}\right\}\subseteq A.$
So, $\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}:(b,x_{l+1},x_{l+2},\ldots,x_{m})\in A\}\in\bigotimes_{i=l+1}^{m}p$ as required.
Now we are going to prove Theorem 1.25.
Proof of Theorem 1.25:.
Similar to the previous discussion we need to prove only $(1)\implies(2).$ This proof is very similar to the above one. We need one technical change in the above proof:
for each $j\in\{1,\ldots,m\}$, let $\alpha_{j}:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ and $H_{j}:\mathbb{P}_{f}(\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ with $\max H_{j}(F)<\min H_{j}(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$; define
$Q(\alpha_{j},H_{j})=\bigcap_{G\in\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})}\overline{T_{G}(\alpha_{j},H_{j})}.$
Now proceed inductively as before. In the each step of the induction choose $p_{j}\in K(Q(\alpha_{j},H_{j}))\subseteq(\mathcal{J},+).$
Now we will prove Theorem 1.27, but before that let us recall one fact about ultrafilters from [14]. We will apply this result. For any discrete semigroup $(S,\cdot)$, and $p\in E\in\left(\beta S,\cdot\right)$, if $A\in p=p\cdot p$, then from [14, Lemma 4.14], the set $A^{\star}=\{x\in A:-x+A\in p\}\in p.$
@@
  2. (22)
    If $n\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\ldots\subsetneq G_{n}\subsetneq K$, then for each $j\in\{1,2,\ldots,m\}$, and $\beta_{1}<\beta_{2}<\dots<\beta_{j}\leq\{n\}$, $P_{1},\ldots,P_{j}\in T,$ and
    $\hskip-199.16928pt\sum_{i\in\beta_{j}}\alpha(G_{i})+P_{j}\left(\sum_{i\in\beta_{j}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\in$
    $D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}.$
@@
Then
$M_{j}\subseteq$
$\bigcap_{\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T}D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}.$
@@
$E=\bigcap^{m}_{j=1}\left\{D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}\right)\right)^{\star}:\right.\\ \left.\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P1,\ldots,P_{j-1}\in T,\text{ and for each }i\in\{1,\dots,n\},f_{i}\in G_{i}\right\}$
$G=\bigcap_{\underset{j=1,\dots,m}{x\in M;}}-x+D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}\in p.$
@@
Now we are going to prove Theorem 1.29.
Proof of Theorem 1.29:.
The proof is similar to the proof of Theorem 1.27.
Let $p_{1},\ldots,p_{m}\in E(K(\beta\mathbb{N},+))$, and $q=\bigotimes_{j=1}^{m}p_{j}$, and let $A\in q$. By using Lifting lemma 1.7, pick for each $j\in\{1,\ldots,m\}$, some $D_{j}:\mathbb{N}^{j-1}\rightarrow\mathcal{P}(\mathbb{N})$ such that
  1. (11)
    for each $j\in\{1,2,\ldots,m\}$, and for each $s\in\{1,2,\ldots,j-1\},w_{s}\in D_{s}(w_{1},w_{2},\dots,w_{s-1})$, and then $D_{j}(w_{1},\dots,w_{j-1})\in p_{j}$, and
  2. (22)
Now we proceed using the induction argument.
For the base case, let $K=\{f\}$. As $D_{1}(\emptyset)^{\star}\in p_{1}$, there exists $a_{1}\in\mathbb{N}$, and $\gamma_{1}\in\mathcal{P}_{f}(\mathbb{N})$ such that for all $P_{1}\in T,a_{1}+P_{1}\left(\sum_{t\in\gamma_{1}}f(t)\right)\in D_{1}(\emptyset)^{\star}$. Define $\alpha_{1}(K)=a_{1}$, and $H_{1}(K)=\gamma_{1}$. Now $\bigcap_{P_{1}\in T}D_{2}\left(a_{1}+P_{1}\left(\sum_{t\in\gamma_{1}}f(t)\right)\right)^{\star}\in p_{2}.$ Again for $K=\{f\}$, there exists $a_{2}\in\mathbb{N}$, and $\gamma_{2}\in\mathcal{P}_{f}(\mathbb{N})$ such that for all $P_{2}\in T,$ we have $a_{2}+P_{2}\left(\sum_{t\in\gamma_{2}}f(t)\right)\in\bigcap_{P_{1}\in T}D_{2}\left(a_{1}+P_{1}\left(\sum_{t\in\gamma_{1}}f(t)\right)\right)^{\star}.$ Define $\alpha_{2}(K)=a_{2}$, and $H_{2}(K)=\gamma_{2}$. Inductively for each $j\in\{1,\ldots,m\},$ we have
  1. (11)
    $a_{j}\in\mathbb{N}$, and $\gamma_{j}\in\mathcal{P}_{f}(\mathbb{N})$ such that for all $P_{j}\in T,$
    $a_{j}+P_{j}\left(\sum_{t\in\gamma_{2}}f(t)\right)\in\bigcap_{P_{1},\ldots,P_{j}\in T}D_{j}\left(a_{1}+P_{1}\left(\sum_{t\in\gamma_{1}}f(t)\right),\ldots,a_{j-1}+P_{j-1}\left(\sum_{t\in\gamma_{j-1}}f(t)\right)\right)^{\star}\in p_{j},$
    and
  2. (22)
    $\alpha_{j}(K)=a_{j}$, and $H_{j}(K)=\gamma_{j}$.
Now assume that $|K|>1$, and that for each $j\in\{1,\ldots,m\},$ $\alpha_{j}(G)$, and $H_{j}(G)$ have been defined for all non-empty proper subsets $G$ of $K$ satisfying
  1. (11)
  2. (22)
    If $n\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\ldots\subsetneq G_{n}\subsetneq K$, then for each $j\in\{1,2,\ldots,m\}$, and $\beta_{1}<\beta_{2}<\dots<\beta_{j}\leq\{n\}$, $P_{1},\ldots,P_{j}\in T,$ and
    $\hskip-199.16928pt\sum_{i\in\beta_{j}}\alpha_{j}(G_{i})+P_{j}\left(\sum_{i\in\beta_{j}}\sum_{t\in H_{j}(G_{i})}f_{i}(t)\right)\in$
    $D_{j}\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha_{j-1}(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H_{j-1}(G_{i})}f_{i}(t)\right)\right)^{\star}.$
for each $j\in\{1,\ldots,m\},$ let $R_{j}=\left\{\sum_{i\in\beta_{j}}\sum_{t\in H_{j}(G_{i})}f_{i}(t):\beta_{j}\leq\{n\},f_{i}\in G_{i}\right\}$, $F=\bigcup\{H_{j}(G):\emptyset\neq G\subsetneq K\text{ and }j\in\{1,\ldots,m\}\}$. Let $M=\max F$ and, let for each $j\in\{1,2,\ldots,m\}$,
$M_{j}=\left\{\sum_{i\in\beta_{j}}\alpha_{j}(G_{i})+P\left(\sum_{i\in\beta_{j}}\sum_{t\in H_{j}(G_{i})}f_{i}(t)\right):\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T,\text{ and for each }i\in\{1,\ldots,n\};f_{i}\in G_{i}\right\}.$
Then
$M_{j}\subseteq$
$\bigcap_{\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T}D_{j}\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha_{j-1}(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H_{j-1}(G_{i})}f_{i}(t)\right)\right)^{\star}.$
Define $R=\bigcup_{j}R_{j}.$ Let $P\in T,d\in R$, define the polynomial $Q_{p,d}$ as $Q_{p,d}(y)=P(y+d)-P(d)$. Clearly $Q_{p,d}\in\mathbb{P}$. Let $S=T\bigcup\{Q_{p,d}:P\in T\text{ and }d\in R\}$.
For each $j\in\{1,\ldots,m\},$ define
$E_{j}=\left\{D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}\right)\right)^{\star}:\right.\\ \left.\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P_{1},\ldots,P_{j-1}\in T,\text{ and for each }i\in\{1,\dots,n\},f_{i}\in G_{i}\right\}$
$G_{j}=\bigcap_{x\in M_{j}}-x+D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P_{j-1}\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}\in p_{j}.$
Now by [7, Lemma 10], pick $b_{1}\in\mathbb{N}$, and $\delta_{1}\in\mathcal{P}_{f}(\mathbb{N})$ such that $\min\delta_{1}>m$, and for all $Q\in S$, and all $f\in K$,
$b_{1}+Q\left(\sum_{t\in\delta_{1}}f(t)\right)\in G_{1}.$
Let $\alpha_{1}(K)=b_{1}$, and $H_{1}(K)=\delta_{1}$.
Now $E_{1}\cap G_{2}\cap D_{1}\left(\left\{b_{1}+Q\left(\sum_{t\in\delta_{1}}f(t)\right):f\in K,Q\in S\right\}\right)^{\star}\in p_{2}.$ Now again by [7, Lemma 10], pick $b_{2}\in\mathbb{N}$, and $\delta_{2}\in\mathcal{P}_{f}(\mathbb{N})$ such that $\min\delta_{2}>m$, and for all $Q\in S$, and all $f\in K$,
$b_{2}+Q\left(\sum_{t\in\delta_{2}}f(t)\right)\in E_{1}\bigcap G_{2}\bigcap D_{1}\left(\left\{b_{1}+Q\left(\sum_{t\in\delta_{1}}f(t)\right):f\in K,Q\in S\right\}\right)^{\star}.$
Define $\alpha_{2}(K)=b_{2}$, and $H_{2}(K)=\delta_{2}$. Now for each $j\in\{1,\ldots,m\}$, inductively one can construct $\alpha_{j}(K)=b_{j}$, and $H_{j}(K)=\delta_{j}$ such that
$b_{j}+Q\left(\sum_{t\in\delta_{j}}f(t)\right)\in E_{j-1}\bigcap G_{j}\bigcap D_{j-1}\left(\left\{\left(b_{1}+Q\left(\sum_{t\in\delta_{1}}f(t)\right)\ldots,b_{j-1}+Q\left(\sum_{t\in\delta_{j-1}}f(t)\right)\right):f\in K,Q\in S\right\}\right)^{\star}.$
Now the verification of the induction hypothesis is similar to the last part of the proof of Theorem 1.27. So we omit the rest.
Now we show that ultrafilters witnessing the conclusion of Theorem 1.29 forms a left ideal of $\beta((\mathbb{N}^{m}),+)$ and $\beta((\mathbb{N}^{m}),\cdot).$ Unfortunately we can’t prove the analogous result for Theorem 1.27. Define
[…]
The following corollary follows from Theorem 1.29.
Corollary 2.1.
$\mathcal{L}$ is a left ideal of $(\beta(\mathbb{N}^{m}),\cdot),$ and $(\beta(\mathbb{N}^{m}),+).$
Proof.
Let $p\in\mathcal{L}$, and $q\in\beta(\mathbb{N}^{k})$ and $B\in q\cdot p.$ Let $T\in\mathcal{P}_{f}(P)$ be given. Then there exists $(x_{1},\ldots,x_{m})\in\mathbb{N}^{m}$ such that $A=(x_{1},\ldots,x_{m})^{-1}B\in p.$ Consider the new set of polynomials $T^{\prime}=\{\frac{1}{x}P:P\in T,\text{ and }x\in\{x_{1},\ldots,x_{m}\}\}.$ Now from Theorem 1.29 , there exist functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H_{,}\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then for every $i\in\{1,\ldots,m\}$, we have $H_{i}(G)<H_{i}(K)$, and
  2. 2.
    if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+\frac{1}{x_{1}}P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+\frac{1}{x_{m}}P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in A.$
For each $i\in\{1,\ldots,m\},$ define $\alpha^{\prime}_{i}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ as $\alpha^{\prime}_{i}=x_{i}\cdot\alpha_{i}$. Then the new set of functions $(\alpha^{\prime}_{i})_{i=1}^{m}$, and $(H_{i})_{i=1}^{m}$ works for the set $B.$ Hence $\overline{B}\cap\mathcal{L}\neq\emptyset.$ This completes the proof.
The proof of $\mathcal{L}$ is a left ideal of $(\beta(\mathbb{N}^{m}),+)$ is easy to derive so we omit the proof. ∎
@@
The following theorem shows that the combinatorial consequence of our Corollary 1.28 is equivalent to a version of the polynomial extension of [14, Theorem 17.31]. The proof uses simple color induction arguments.
@@
    Let $r,m\in\mathbb{N}$, and let $\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$. Then there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
      $\left\{\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right):\right$.
      $\hskip 289.07999ptf_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\Bigg\}\subset C_{j}$
@@
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have for every $f_{i}\in G_{i},$ and $P_{1},\ldots,P_{m}\in T,$
      $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in D_{j}$